A single stage reciprocating compressor takes 1 m3 of air per minute at 1.013 bar and 15°C and delivers it at 7 bar. Assuming that the law of compression ispv1.35=c , and that clearance volume is negligible, calculate the indicated power.



1
Expert's answer
2021-07-09T08:31:37-0400

"p_1=1.013 \;bar \\

T_1=288 \;K \\

V_1= 1 \;m^3/min \\

p_2= 7 \;bar \\

p_v^{1.35}=c \\

m_a= \frac{p_1V_1}{RT_1} \\

= \frac{1.013 \times 10^2 \times 1/60}{0.287 \times 288} = 0.02042 \;kg/s"

Delivery temperature

"\frac{T_2}{T_1}=(\frac{p_2}{p_1})^{\frac{n-1}{n}} \\

T_2=288(\frac{7}{1.013})^{\frac{0.35}{1.35}} = 475.2\;K"

The indicated power "= \frac{n}{n-1}mR(T_2-T_1)"

"= \frac{1.35}{0.35} \times 0.02042 \times 0.287 \times (475.2-288) =4.23 \;kW"